40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 12 16:57:38 2011, Jenne, Update, RF System, RAMmon, 4 day trend EOM_tempSense_noFoam_noHeater_RAMmon_4days.png
    Reply  Tue Dec 13 11:34:32 2011, Jenne, Update, RF System, RAMmon, 5 day trend EOM_tempSense_RAMmon_5days_last_day_withFoam_noHeater.pngEOM_tempSense_withFoam_noHeater_RAMmon_1day.png
Message ID: 6111     Entry time: Tue Dec 13 11:34:32 2011     In reply to: 6109
Author: Jenne 
Type: Update 
Category: RF System 
Subject: RAMmon, 5 day trend 

Now we've got another day of data, with the foam box on for the last 24hrs.

First plot is a 5 day trend so you can see that the RAM has gotten a little bit smaller, as has the temperature drift, but not by a whole lot.

EOM_tempSense_RAMmon_5days_last_day_withFoam_noHeater.png

Second plot is the last 19 hours (so excluding much of the time while I was realigning beams on the PSL table yesterday), to zoom in on just the time when the foam box was installed.

EOM_tempSense_withFoam_noHeater_RAMmon_1day.png

After lunch Zach and I are going to engage the heater to temperature stabilize the system, to see how that affects things.

In other news, the MC looks like it was fine for a good long time, and ~3 hours ago it went bad.  The mode that's flashing is really bad in both pitch and yaw.  I don't know what happened, but something is not so awesome. Edit: Steve said that he opened the PSL table at some point this morning to look around but not touch, and also it's Janitor Day, and Kevin comes in around 8ish. That doesn't mean I know the actual cause, but those are the only things that happened in the IFO room this morning that anyone is aware of.

ELOG V3.1.3-